If t n be the n th term of an AP and if t 7 =9, then the value of common difference that would make t 1 t 2 t 7 least is____: A)33/40 B)33/20 C)33/10

If tn be the nth term of an AP and if t7=9, then the value of common difference that would make t1t2t7 least is____:
A)33/40     B)33/20    C)33/10

Grade:11

1 Answers

Arun
6 years ago
Dear Shiv
 
Since 
T1 T2 T7/9
= (9 -6d) (9-5d)
= 30 d² -99d +81
= 30 (d²- 99d/30) + 81
= 30 (d² - 33d/10 + 33²/20²) + 81 - 30* 33²/20²
= 30 ( d-33/20)² + 270/400
 
To make it least
 
d - 33/20 = 0
Hence
d = 33/20
 
 
Regards
Arun (askIITians forum expert)